### v1

# New axisymmetric equilibria with flow from

Abstract
We construct analytic solutions to the generalized Grad-Shafranov (GS) equation, adopting the general linearizing ansatz for the free-function terms it contains, by expanding the genaralized Solovév solution [Ch. Simintzis, G. N. Throumoulopoulos, G. Pantis and H. Tasso, Phys. Plasmas 8, 2641 (2001)]. On the basis of these solutions, we examine how the genaralized Solovév configuration is modified as the values of the free parameters associated with the additional pressure, poloidal-current and electric-field terms are changed. Thus, a variety of equilibria of tokamak, spherical tokamak and spheromak pertinence are constructed including D-shaped configurations with positive and negative triangularity and diverted configurations with either a couple of X-points or a single X-point.
1 Introduction
The magnetohydrodynamics axisymmetric equilibrium states are governed by the well-known GS equation, a quasi-linear elliptic partial differential equation for the poloidal magnetic flux-function, containing a couple of arbitrary functions of that flux (surface functions), namely the pressure and the poloidal-current functions. This equation is generally solved numerically under appropriate boundary conditions, and to this end, several codes have been developed, e.g. the HELENA code [1]. Also, to easier gain physical intuition and for benchmarking equilibrium codes, analytic solutions have been constructed to linearized forms of the GS equation. In connection with the present study we first mention the widely employed in plasma confinement studies Solovév solution [2]. This solution, corresponding to constant surface-function terms in the GS equation, describes an up-down symmetric equilibrium with D-shaped magnetic surfaces surrounded by a spontaneously formed separatrix with a couple of X-points (cf. Fig. 1). Also, known is the Maschke-Hernegger solution [3, 4], expressed in terms of the Whittaker functions, corresponding to linear choices of the surface-function terms and describing more realistic equilibria with current-density profiles vanishing on the plasma boundary. Several additional analytic solutions to the GS equation are available, e.g. extending the Solovév solution by polynomial contributions to the homogeneous counterpart of the GS equation in order to construct diverted tokamak equilibria [5]; for constant pressure and linear current-density choices of the free terms in the GS equation [6]; for linear choices of both the pressure and current-density terms by imposing D-shaped boundaries or boundaries with a lower X-point [7, 8, 9]; by differentiation of known separated solutions with respect to the separation parameter in order to get simpler solutions [10]; and in elliptic prolate geometry [11]. Also, for the case of the generic linearized form of the GS equation an analytic solution involving infinite series was derived in [12] and was employed to construct equilibrium configurations pertinent to the tokamak ASDEX-Upgrade.
Extended equilibrium investigations have been conducted to include plasma flows and associated electric fields, which play an important role in the creation of advanced confinement regimes in tokamaks, e.g. the transition of the low to high confinement mode. In particular, for incompressible flows, the generalized GS equation (1) shown below was derived [13, 14], containing five free surface quantities. Equation (1) has an additional electric-field term through the electrostatic potential $\Phi$, associated with the component of plasma velocity non-parallel to the magnetic field. For parallel velocity and arbitrary assignment of the free surface-function terms, an extension of the HELENA code was developed in [15], while the generalized Solovév solution to (1) (given by (5) below), which constitutes a basic ingredient of the present study, was obtained in [14]. Owing to the electric field, the respective equilibrium in addition to a separatrix, similar to that of the usual static Solovév equilibrium, possess a third X-point outside the separatrix (cf. Fig. 11 of [14]). Other analytic solutions to the generalized GS equation of tokamak or space plasma pertinence were obtained by adopting a Solovév-like linearizing ansatz and imposing a diverted boundary [16]; alternative linearizing ansatzes [17, 18, 19]; non-linear forms of the generalized GS equation [20]; employing Lie-point symmetries for linear and nonlinear choices of the free-function terms [21, 22, 23]; and by the method of similarity reduction [24, 25].
Aim of the present work is to construct analytically solutions to the generalized GS equation in its generic linearized form by employing an alternative method. This consists in pursuing solutions as expanded generalized Solovév ones, i.e. as superposition of the generalized Solovév solution and a function to be determined. In addition to the novelty of this method we are particularly interested in examining how the Solovév configuration is modified when the influence of the expanding function becomes stronger by changing the pertinent new free parameters.
In section 2 the generalized GS equation and the generalized Solovév solution are briefly reviewed and the method of solving the generalized GS equation adopting the generic linearized ansatz is presented. Then, analytic solutions to the generic linearized generalized GS equation are constructed in section 3. In section 4 we examine how the Solovév configuration is modified by varying the additional free parameters associated with the pressure and the poloidal current in conjunction with the parallel component of the velocity and electric field, thus obtaining a variety of new configurations. The conclusions are summarized in section 5.
2 Generic linear axisymmetric equilibria with non-parallel flow
The generalized GS equation is given in SI units by (cf. [13, 14])
$\Delta^{*}u+\frac{1}{2}\frac{d}{du}\left[\frac{X^{2}}{1-M_{p}^{2}}\right]+\mu_{0}R^{2}\frac{dP_{s}}{du}+\mu_{0}\frac{R^{4}}{2}\frac{d}{du}\left[\rho\left(\frac{d\Phi}{du}\right)^{2}\right]=0\,.$ (1)
Here $(R,\phi,z)$ are cylindrical coordinate with $z$ corresponding to the axis of symmetry; the poloidal magnetic flux function $u(R,z)$ labels the magnetic surfaces; $M_{p}(u)$ is the Mach function of the poloidal velocity with respect to the poloidal Alfven velocity; $\Phi(u)$ is the electrostatic potential; $\rho(u)$ is the plasma density; $X(u)$ relates to the toroidal magnetic field $B_{\phi}=I/R$ through
$I=\frac{X}{\left(1-M_{p}\right)}-R^{2}\frac{\sqrt{\rho}M_{p}\Phi^{\prime}}{\left(1-M_{p}\right)^{1/2}}\,.$ (2)
The operator $\Delta^{*}$ in (1) is $\Delta^{*}:=\partial^{2}/\partial R^{2}-(1/R)\partial/\partial R+\partial^{2}/\partial z^{2}$ and $\mu_{0}$ is the permeability of free space. For vanishing flow, i.e. $M_{p}=\Phi^{\prime}=0$, the surface function $P_{s}(u)$ coincides with the pressure and Eq. (1) reduces to the GS equation.
Following Ref. [14] we make the following choice of the free surface-function terms in (1)
$\frac{d(\mu_{0}P_{s})}{du}$ $=$ $-\frac{P_{s0}}{U_{0}}\,,$
$\frac{1}{2}\frac{d}{du}\left[\frac{\mu_{0}X^{2}}{1-M_{p}^{2}}\right]$ $=$ $\epsilon\frac{P_{s0}R_{0}^{2}}{U_{0}(1+\delta^{2})}\,,$
$\frac{1}{2}\frac{d}{du}\left[\mu_{0}\rho\left(\frac{d\Phi}{du}\right)^{2}\right]$ $=$ $-\lambda\frac{P_{s0}}{U_{0}(1+\delta^{2})R_{0}^{2}}\,.$ (3)
Normalizing as $x:=R/R_{0}$, $y:=z/R_{0}$, $U:=u/u_{0}$, $u_{0}:=P_{s0}R_{0}^{4}/(2U_{0}(1+\delta^{2}))$, Eq. (1) assumes the form
$U_{xx}-\frac{1}{x}U_{x}+U_{yy}-2\epsilon x^{4}+-2(1+\delta^{2})x^{2}+2\epsilon=0\,,$ (4)
and the generalized Solovév solution is given by
$U_{gs}=\left[y^{2}\left(x^{2}-\epsilon\right)+\frac{\delta^{2}+\lambda}{4}(x^{2}-1)^{2}+\frac{\lambda}{12}(x^{2}-1)^{3}\right]\,.$ (5)
The equilibrium configuration for $\lambda>0$ is shown in Fig. 1.
Figure 1: The equilibrium configuration of the GS solution (5). The red-dashed curve represents the separatrix with $x_{s1}=\sqrt{\epsilon}$, $x_{s2}=\left(-\epsilon\lambda-3\delta^{2}+\sqrt{3}\sqrt{-\epsilon^{2}\lambda^{2}-2\delta^{2}\epsilon\lambda+3\delta^{4}+\delta^{2}\lambda+4\lambda^{2}}/(2\lambda)\right)^{1/2}$ and $y_{s}=\left[(1-\epsilon)^{2}(2\delta^{2}+\lambda(\epsilon+1))\right]^{1/2}$.
As already mentioned in the Introduction, the outermost closed surface (red-dashed line) is a separatrix consisting of an elliptic outer part and a straight inner part parallel to the axis of symmetry; thus, a couple of X-points are created. Expressions of the X-point coordinates and of the inner and outer points of the configuration on the mid-plane $y=0$ in terms of the parameters $\delta$, $\epsilon$ and $\lambda$ are given in the caption of Fig. 1. The magnetic axis is located at ($x=1,y=0$). The parameter $\epsilon$ determines the compactness of the configuration, i.e. for $\epsilon>0$ the equilibrium is diamagnetic and describes either a tokamak or a spherical tokamak, while for $\epsilon=0$ the inner part of the separatrix touches the axis of symmetry forming a spheromak. The parameter $\delta$ relates to the elongation of the magnetic surfaces in the vicinity of the magnetic axis, i.e. the higher the value of $\delta$, the larger the elongation parallel to the axis of symmetry. For $\lambda=0$ the velocity becomes parallel to the magnetic field, while for $\lambda<0$ a third X-point appears outside the separatrix. When the velocity vanishes ($M_{p}=\lambda=0$), the usual static Solovév solution is recovered. Here we will consider equilibria with $\lambda>0$.
We now try the following new generic linearizing ansatz to Eq. (1)
$\frac{d(\mu_{0}P_{s})}{du}$ $=$ $-\frac{P_{s0}}{U_{0}}+\bar{\alpha}\frac{P_{s0}}{U_{0}^{2}}u\,,$
$\frac{1}{2}\frac{d}{du}\left[\frac{\mu_{0}X^{2}}{1-M_{p}^{2}}\right]$ $=$ $\epsilon\frac{P_{s0}R_{0}^{2}}{U_{0}(1+\delta^{2})}-\bar{\beta}\epsilon\frac{P_{s0}R_{0}^{2}}{U_{0}^{2}(1+\delta^{2})}u\,,$
$\frac{1}{2}\frac{d}{du}\left[\mu_{0}\rho\left(\frac{d\Phi}{du}\right)^{2}\right]$ $=$ $-\lambda\frac{P_{s0}}{U_{0}(1+\delta^{2})R_{0}^{2}}+\bar{\gamma}\frac{\lambda P_{s0}}{U_{0}^{2}(1+\delta^{2})R_{0}^{2}}u\,,$ (6)
where $\bar{\alpha},\bar{\beta},\bar{\gamma}$ are dimensionless parameters. Setting $G:=2\epsilon$, $F:=-2(1+\delta^{2})$, $E:=-2\lambda$, $\alpha:=\bar{\alpha}(u_{0}/U_{0})$, $\beta:=\bar{\beta}(u_{0}/U_{0})$ and $\gamma:=\bar{\gamma}(u_{0}/U_{0})$ we obtain from (1) the following equation:
$U_{xx}-\frac{1}{x}U_{x}+U_{yy}+Ex^{4}+Fx^{2}+G-Ex^{4}\gamma U-Fx^{2}\alpha U-G\beta U=0\,.$ (7)
Note that the solution to (7) for $\alpha=\beta=\gamma=0$ is given by (5). We now expand about the GS solution (5) as
$U=U_{gs}+W(x,y)\,,$ (8)
with the function $W(x,y)$ to be determined so that (8) to satisfy (7). Substituting (8) into (7) yields
$W_{xx}-\frac{1}{x}W_{x}+W_{yy}-\gamma Ex^{4}W-\alpha Fx^{2}W-\beta GW=$
$=(\gamma Ex^{4}+\alpha Fx^{2}+\beta G)U_{gs}\,.$ (9)
This is an inhomogeneous linear second-order partial differential equation for the unknown function $W(x,y)$. The general solution consists of the sum of the general solution to the counterpart homogeneous equation plus a particular solution of the complete inhomogeneous equation.
3 Solutions
We first consider the homogeneous counterpart of (2):
$W_{xx}-\frac{1}{x}W_{x}+W_{yy}-(\gamma Ex^{4}+\alpha Fx^{2}+\beta G)W=0\,.$ (10)
The general solution of (10) can be found by separation of variables, i.e. in the form
$W(x,y)=X(x)Y(y)\,.$ (11)
Inserting (11) into (10) we find:
$Y^{{}^{\prime\prime}}+k^{2}Y=0\,,$ (12)
and
$X^{{}^{\prime\prime}}-\frac{1}{x}X^{{}^{\prime}}-(\gamma Ex^{4}+\alpha Fx^{2}+\beta G+k^{2})X=0\,,$ (13)
where $k$ is the parameter of separation. The general solution of (12) is
$Y(y)=D_{1}\cos(ky)+D_{2}\sin(ky)\,.$ (14)
Setting $\xi=x^{2}$, Eq. (13) takes the form
$4\xi X^{{}^{\prime\prime}}(\xi)-(\gamma E\xi^{2}+\alpha F\xi+\beta G+k^{2})X=0\,.$ (15)
Since $\xi=0$ is a regular singular point of (15) its solution can be derived by the method of Frobenius in terms of infinite converging series of the form:
$X(\xi)=\sum_{n=0}^{\infty}a_{n}\xi^{n+r}\,.$ (16)
Substituting (16) into (15) results in the following recursion relations:
$4r(r-1)a_{0}=0\,,$
$4r(r+1)a_{1}-(\beta G+k^{2})a_{0}=0\,,$
$4(r+1)(r+2)a_{2}-\alpha Fa_{0}-(\beta G+k^{2})a_{1}=0\,,$
[…]
where $a_{0}$ is a freely specified parameter. The roots of the indicial polynomial, $r(r-1)$, are $r_{1}=1$ and $r_{2}=0$. Therefore, for $r_{1}=1$, one solution of (15) is
$X_{1}(\xi)=\xi\sum_{n=0}^{\infty}a_{n}\xi^{n}\,.$ (18)
Since $r_{1}-r_{2}$ is a non-zero integer, the second linearly independent solution is of the form
$X_{2}(\xi)=cX_{1}(\xi)\ln\xi+\sum_{n=0}^{\infty}\bar{a}_{n}\xi^{n}\,,$ (19)
where the coefficients $c$ and $\bar{a}_{n}$ can be determined by substituting (19) into (15). Here, we will employ the solution $X_{1}(\xi)$.
Furthermore, we will pursue a particular solution of (2) in the form:
$W(x,y)=W_{1}(x)y^{2}+W_{2}(x)\,.$ (20)
Inserting (20) into (2) yields
$W_{1}^{{}^{\prime\prime}}-\frac{1}{x}W_{1}^{{}^{\prime}}-(\gamma Ex^{4}+\alpha Fx^{2}+\beta G)W_{1}=\left(x^{2}-\frac{G}{2}\right)(\gamma Ex^{4}+\alpha Fx^{2}+\beta G)\,,$ (21)
and
$W_{2}^{{}^{\prime\prime}}-\frac{1}{x}W_{2}^{{}^{\prime}}+2W_{1}-(\gamma Ex^{4}+\alpha Fx^{2}+\beta G)W_{2}=$
$=-\left[\frac{(F+E+2)}{8}(x^{2}-1)^{2}+\frac{E}{24}(x^{2}-1)^{3}\right](\gamma Ex^{4}+\alpha Fx^{2}+\beta G)\,.$ (22)
Equation (21) involves only the function $W_{1}(x,y)$. Thus, once (21) is solved, the solution contributes to the inhomogeneous part of (3).
To solve (21) we again change variable, $\xi=x^{2}$, to get:
$4\xi W_{1}^{{}^{\prime\prime}}(\xi)-(\gamma E\xi^{2}+\alpha F\xi+\beta G)W_{1}=\left(\xi-\frac{G}{2}\right)(\gamma E\xi^{2}+\alpha F\xi+\beta G)\,.$ (23)
Setting
$W_{1}+\left(\xi-\frac{G}{2}\right):=\bar{W}_{1}\,,$ (24)
Eq. (23) is put in the form:
$4\xi\bar{W}_{1}^{\prime\prime}-(\gamma E\xi^{2}+\alpha F\xi+\beta G)\bar{W}_{1}=0\,.$ (25)
For $\beta=0$, the solution of (25) is:
$\bar{W}_{1}=c_{1}Ai\left(\frac{\alpha F+\gamma E\xi}{(2\gamma E)^{2/3}}\right)+c_{2}Bi\left(\frac{\alpha F+\gamma E\xi}{(2\gamma E)^{2/3}}\right)\,,$ (26)
where $Ai$ and $Bi$ are the Airy functions of the first and second kind. For $\beta\neq 0$ apart from the $k^{2}$-term, Eq. (15) is identical with (25). Therefore a solution of (25) is:
$\bar{W}_{1}(\xi)=\xi\sum_{n=0}^{\infty}d_{n}\xi^{n}\,,$ (27)
with the coefficients $d_{n}$ being determined by the recurrence relations (3) after setting $k=0$ (and replacing $a_{n}$ with $d_{n}$).
Lastly, we will obtain a particular solution of the inhomogeneous equation (3). Introducing once more the variable $\xi=x^{2}$, Eq. (3) is written as:
$4\xi W_{2}^{{}^{\prime\prime}}(\xi)-(\gamma E\xi^{2}+\alpha F\xi+\beta G)W_{2}=$
$=-2W_{1}(\xi)-[C_{5}\xi^{5}+C_{4}\xi^{4}+C_{3}\xi^{3}+C_{2}\xi^{2}+C_{1}\xi+C_{0}]\,,$ (28)
where $W_{1}=\bar{W}_{1}-(\xi-G/2)$ by (24) and
$C_{5}$ $=$ $\gamma\frac{E^{2}}{24}\,,$
$C_{4}$ $=$ $\frac{\gamma E(F+2)}{8}+\alpha F\frac{E}{24}\,,$
$C_{3}$ $=$ $\beta G\frac{E}{24}+\alpha F\frac{(F+2)}{8}-\gamma E\frac{(2F+E+4)}{8}\,,$
$C_{2}$ $=$ $\beta G\frac{(F+2)}{8}-\alpha F\frac{(2F+E+4)}{8}+\gamma E\frac{(3F+2E+6)}{24}\,,$
$C_{1}$ $=$ $-\beta G\frac{(2F+E+4)}{8}+\alpha F\frac{(3F+2E+6)}{24}\,,$
$C_{0}$ $=$ $\beta G\frac{(3F+2E+6)}{24}\,.$ (29)
In view of the converging series solutions (18) and (27) to the respective homogeneous equations (15) and (25), we pursue a similar in form solution to (3):
$W_{2}=\xi\sum_{n=0}^{\infty}b_{n}\xi^{n}\,.$ (30)
Inserting (30) into (3) we obtain after some tedious but straightforward algebra that $b_{0}$ is freely specified and
$C_{0}+G=0\,,$
$8b_{1}-\beta Gb_{0}+2d_{0}-2+C_{1}=0\,,$
$24b_{2}-\alpha Fb_{0}-\beta Gb_{1}+2d_{1}+C_{2}=0\,,$
$48b_{3}-\gamma Eb_{0}-\alpha Fb_{1}-\beta Gb_{2}+2d_{2}+C_{3}=0\,,$
$80b_{4}-\gamma Eb_{1}-\alpha Fb_{2}-\beta Gb_{3}+2d_{3}+C_{4}=0\,,$
$120b_{5}-\gamma Eb_{2}-\alpha Fb_{3}-\beta Gb_{4}+2d_{4}+C_{5}=0\,,$
[…]
Inspection of the last of Eqs. (3) and $C_{0}+G=0$ (first of Eqs. (3)) together with the definitions $G:=2\epsilon$, $F:=-2(1+\delta^{2})$ and $E:=-2\lambda$ implies the following:
  1. 1.
    When $\beta=0$, then $\epsilon=0$ and therefore the second of (2) becomes $(1/2)d/du(X^{2}/(1-M_{p}^{2}))=0$. In the absence of flow then (2) implies that the plasma is confined by a vacuum toroidal magnetic field ($\beta_{p}=1$ equilibrium); when the non-parallel flow is there the electric-field term in (2) additionally contributes to the poloidal current.
  2. 2.
    When $\epsilon\neq 0$ and $\beta\neq 0$, the last of (3) becomes $\beta(3F+2E+6)+24=0$. In this case all the linear terms in the ansatz (2) can remain finite. The latter relation just acts as a constraint on the parameters $\epsilon$, $\delta$ and $\lambda$. To construct the equilibria of Sec. 4 we treated $\epsilon$ and $\lambda$ as free and
    $\delta=\left[\frac{(2/3)(6-\beta\lambda)}{\beta}\right]^{1/2}\,.$ (32)
  3. 3.
    When $\epsilon=0$, $\beta$ does not necessarily vanish and both $\delta$ and $\lambda$ remain free.
Additional free parameters are $\alpha$, $\beta$, $\gamma$, $a_{0}$, $d_{0}$, $b_{0}$, $D_{1},D_{2}$, $R_{0}$, $P_{s0}$, $U_{0}$ and $k$. Also, it is noted that the series expansion (30) was shown numerically to converge and we kept terms up to the order of $\xi^{20}$ in our calculations.
Summarizing, a solution of the general linearized form of the generalized GS equation (2) consists of a separated solution of its homogeneous counterpart in terms of (11), (14) and (18) plus a particular solution of the form (20), expressed in terms of the series (27) and (30).
Figure 2: Left: different equilibrium boundaries are given as the additional linear-in-$U$, pressure term in the first of (2) varies in connection with different values of $\alpha$. The black continuous curve represents the separatrix of the respective genaralized Solovév spheromak configuration. Right: a set of magnetic surfaces for one of the configuration representd by blue, dot-dashed U-curves are shown together with the respective curves of the genaralized Solovév equilibrium.
4 Equilibrium configurations
By employing the solution to (7) derived in the previous section we have constructed several new equilibria. Particularly, we examined how the generalized Solovév equilibrium is modified by varying the additional parameters $\alpha$, $\beta$ and $\gamma$, associated with the respective linear-in-$u$ terms in (2) of pressure, poloidal-current in conjunction with the poloidal velocity, and electric field. To impose a stagnation point, $(x_{0},y_{0})$, in addition to the conditions $dU/dx|_{x_{0},y_{0}}$ and $dU/dy|_{x_{0},y_{0}}$ we considered the sign of the quantity
$\left.\frac{d^{2}U}{dx^{2}}\frac{d^{2}U}{dy^{2}}-\left(\frac{d^{2}U}{dxdy}\right)^{2}\right|_{(x_{0},y_{0})}\,,$
which should be positive for a magnetic axis and negative for an X-point. For the up-down symmetric configurations considered, the magnetic axis is located at the same position, $(x=1,y=0)$ as the generalized Solovév configuration and $U=U_{a}=0$ thereon. Note that in this case the condition $dU/dy|_{(1,0)}$ is identically satisfied because of symmetry. If not otherwise specified, we assigned the following parametric values: $k=0.1$, $\delta=0.1$, $\lambda=0.1$, $D_{1}=1$ and $D_{2}=0$. It is recalled that when $\beta\neq 0$ and $\epsilon\neq 0$, $\delta$ should be calculated from (32). The parameters $a_{0}$, $b_{0}$ and $d_{0}$ were determined as solutions to the set of the algebraic equations derived by the conditions associated with the imposition of a stagnation point. Since we employed dimensionless quantities, the reference dimensional parameters $R_{0}$, $P_{s0}$ (in units of $4\pi\times 10^{-7}$ $H/m\,Pa$ ) and $U_{0}$ remained free. Depending on the value of $\epsilon$ we constructed tokamak, spherical tokamak and spheromak equilibria, the latter corresponding to $\epsilon=0$.
First, we examined the impact of the $\alpha$- $\beta$- and $\gamma$-term individually on up-down symmetric equilibria equilibria by imposing the same position of the magnetic axis, $(x=1,y=0)$, as that of the respective generalized Solovév equilibrium and the same value $U=0$ thereon. In the presence of $\alpha$-term, a D-shaped configuration is formed with negative triangularity, the boundary of which is shown in Fig. 2-left by the red-dashed curve. As $\alpha$ gradually takes larger values the configuration increasing triangularity extends towards the axis of symmetry, touches the axis of symmetry forming a separatrix with a single X-point located on the origin of the coordinate system (blue, dot-dashed line) and then becomes a peculiar configuration having a separatrix with a couple of X-points (green-dotted line).
The $\beta$-term creates configurations with small triangularity elongated along the $x$-axis as shown in Fig. 3. The red-dashed curves represent the magnetic surfaces of a configuration with the same poloidal magnetic flux as the respective generalized Solovév tokamak configuration represented by the black-continuous curve. As $\beta$ takes larger values the configuration gets more extended and eventually reaches the axis of symmetry becoming a spheromak one. It is noted that for $\epsilon=\beta=\gamma=0$ the equilibrium becomes independent of $\beta$. The impact of the $\gamma$-term is similar to that of the $\alpha$-term as shown in Fig. 4.
Subsequently, the impact of combinations of the $\alpha$-, $\beta$- and $\gamma$-terms was examined by imposing an X-point at several positions. Thus, a variety of configurations were created. Three examples are given in Figs. 5-7. In Fig. 5 the blue, dot-dashed curves represent an up-down symmetric spherical tokamak configuration ($\epsilon=0.1$) the X-points of which are common with those of the respective generalized Solovév equilibrium, but the poloidal magnetic flux of this former equilibrium is larger than that of the latter. The red-dashed curve represents the boundary of a D-shaped configuration with poloidal magnetic flux equal to that of the generalized Solovév solution. In Fig. 6 is shown an up-down symmetric diverted configuration with the imposed X-points located at ($x=0.84,y=\pm 14.5$), while Fig. 7 shows an up-down asymmetric configuration with the magnetic axis located at ($x=1.02,y=0.83$) and the single lower X-point at $(x=0.94,y=-22.6)$.
Figure 3: The impact of the additional linear-in-$U$, poloidal-current term in the second of (2) in connection with different values of $\beta$. The black continuous curves represent the magnetic surfaces of the respective generalized Solovév tokamak configuration.
Figure 4: Left: different equilibrium boundaries are given as the additional linear-in-$U$, electric-field term in the third of (2) varies in connection with different values of $\gamma$. The black continuous curve represents the separatrix of the respective generalized Solovév spheromak configuration. Right: A set of magnetic surfaces for one of the configuration represented by green-dotted U-curves are shown together with the respective curves of the Solovév equilibrium.
Figure 5: A spherical tokamak diverted configuration ($\epsilon=0.1$, $\delta=20$) with a couple of X-points common with the respective Solovév equilibrium. The red-dashed curve represents the boundary of a D-shaped tokamak equilibrium with poloidal magnetic flux equal to that of the Solovév one.
Figure 6: An up-down symmetric, diverted tokamak configuration ($\epsilon=0.1,\ \delta=20$) with a couple of X-points located at ($x=0.84,y=\pm 14.5$). The black continuous curves represent the magnetic surfaces of the respective Solovév equilibrium.
Figure 7: An up-down asymmetric, diverted tokamak configuration for $k=0.2$, $\epsilon=0.2$, $\delta=31.6$, $D_{1}=0.5$ and $D_{2}=0.01$. The magnetic axis is located at $(x=1.02,y=0.83)$ and the single lower X-point at $(x=0.94,y=-22.6)$. The black continuous curves represent the magnetic surfaces of the respective Solovév equilibrium.
5 Conclusions
Making a generic linearizing choice of the free function-terms involved in the generalized GS equation (1) we solved the resulting linearized equation (2) analytically. The requested solution was expressed as an expansion of the generalized Solovév solution (5) in the form (8) involving the determinable function $W$. Then, the solution was obtained as a superposition of the counterpart homogeneous equation and a particular solution of the complete inhomogeneous equation in the form (20). The solutions of the pertinent radial ODEs were expressed in terms of infinite converging series.
Employing these solutions, we constructed several configurations by changing the values of the free parameters, either individually or in combination, associated with the new terms of pressure, poloidal current and electric field. The equilbria constructed are pertinent to tokamaks, spherical tokamaks and spheromaks; they include D-shaped configurations with either positive or negative triangularity and diverted configurations with either a couple of X-points or a single X-point.
The present study can be extended to a more generic generalized GS equation accounting for CGL pressure anisotropy ([26], Eq. (34) therein). Another extension could be pursued in the framework of Hall-MHD, a simplified two-fluid model with inertialess electron-fluid elements keeping on magnetic surfaces and ion-fluid elements departing from them.
Aknowledgments
This work was conducted in the framework of participation of the University of Ioannina in the National Programme for the Controlled Thermonuclear Fusion, Hellenic Republic.

### v2

@@
 Abstract
-We construct analytic solutions to the generalized Grad-Shafranov (GS) equation, adopting the general linearizing ansatz for the free-function terms it contains, by expanding the genaralized Solovév solution [Ch.
+We construct analytic solutions to the generalized Grad-Shafranov equation, which incorporates both toroidal and poloidal flows.
+This is achieved by adopting a general linearizing ansatz for the free-function terms of the equation and expanding the generalized Solov’ev solution [Ch.
 Simintzis, G.
@@
 Plasmas 8, 2641 (2001)].
-On the basis of these solutions, we examine how the genaralized Solovév configuration is modified as the values of the free parameters associated with the additional pressure, poloidal-current and electric-field terms are changed.
-Thus, a variety of equilibria of tokamak, spherical tokamak and spheromak pertinence are constructed including D-shaped configurations with positive and negative triangularity and diverted configurations with either a couple of X-points or a single X-point.
+On the basis of these solutions, we examine how the genaralized Solov’ev configuration is modified as the values of the free parameters associated with the additional pressure, poloidal-current and electric-field terms are changed.
+Thus, a variety of equilibria of tokamak, spherical tokamak and spheromak pertinence are constructed, including D-shaped configurations with positive and negative triangularity and diverted configurations with either a couple of X-points or a single X-point.
 1 Introduction
-The magnetohydrodynamics axisymmetric equilibrium states are governed by the well-known GS equation, a quasi-linear elliptic partial differential equation for the poloidal magnetic flux-function, containing a couple of arbitrary functions of that flux (surface functions), namely the pressure and the poloidal-current functions.
+The magnetohydrodynamics axisymmetric equilibrium states are governed by the well-known Grad-Shafranov (GS) equation, a quasi-linear elliptic partial differential equation for the poloidal magnetic flux-function, containing a couple of arbitrary functions of that flux (surface functions), namely the pressure and the poloidal-current functions.
 This equation is generally solved numerically under appropriate boundary conditions, and to this end, several codes have been developed, e.g.
@@
 Also, to easier gain physical intuition and for benchmarking equilibrium codes, analytic solutions have been constructed to linearized forms of the GS equation.
-In connection with the present study we first mention the widely employed in plasma confinement studies Solovév solution [2].
+In connection with the present study we first mention the widely employed in plasma confinement studies Solov’ev solution [2].
 This solution, corresponding to constant surface-function terms in the GS equation, describes an up-down symmetric equilibrium with D-shaped magnetic surfaces surrounded by a spontaneously formed separatrix with a couple of X-points (cf.
@@
 Several additional analytic solutions to the GS equation are available, e.g.
-extending the Solovév solution by polynomial contributions to the homogeneous counterpart of the GS equation in order to construct diverted tokamak equilibria [5]; for constant pressure and linear current-density choices of the free terms in the GS equation [6]; for linear choices of both the pressure and current-density terms by imposing D-shaped boundaries or boundaries with a lower X-point [7, 8, 9]; by differentiation of known separated solutions with respect to the separation parameter in order to get simpler solutions [10]; and in elliptic prolate geometry [11].
+extending the Solov’ev solution by polynomial contributions to the homogeneous counterpart of the GS equation in order to construct diverted tokamak equilibria [5]; for constant pressure and linear current-density choices of the free terms in the GS equation [6]; for linear choices of both the pressure and current-density terms by imposing D-shaped boundaries or boundaries with a lower X-point [7, 8, 9]; by differentiation of known separated solutions with respect to the separation parameter in order to get simpler solutions [10]; and in elliptic prolate geometry [11].
 Also, for the case of the generic linearized form of the GS equation an analytic solution involving infinite series was derived in [12] and was employed to construct equilibrium configurations pertinent to the tokamak ASDEX-Upgrade.
@@
 the transition of the low to high confinement mode.
-In particular, for incompressible flows, the generalized GS equation (1) shown below was derived [13, 14], containing five free surface quantities.
-Equation (1) has an additional electric-field term through the electrostatic potential $\Phi$, associated with the component of plasma velocity non-parallel to the magnetic field.
-For parallel velocity and arbitrary assignment of the free surface-function terms, an extension of the HELENA code was developed in [15], while the generalized Solovév solution to (1) (given by (5) below), which constitutes a basic ingredient of the present study, was obtained in [14].
-Owing to the electric field, the respective equilibrium in addition to a separatrix, similar to that of the usual static Solovév equilibrium, possess a third X-point outside the separatrix (cf.
+Specifically, to include axisymmetric toroidal flows, inherently incompressible because of the symmetry, elliptic PDEs were derived in [13] for either isentropic or isothermal magnetic surfaces and analytic solutions were constructed therein.
+For compressible flows of arbitrary direction the problem reduces to a set of a PDE coupled with a Bernoulli equation involving the pressure [14, 15, 16].
+Depending on the value of the poloidal velocity, this PDE can be either elliptic or hyperbolic, that is, there are three critical transition poloidal velocities, thus forming two elliptic and two hyperbolic regions.
+In the frame of two-fluid model, the electron and ion surfaces depart from the magnetic surfaces and the problem reduces to a set of three PDEs involving respective flux-function labels [17, 18].
+The problem was also investigated within the framework of simplified two-fluid models, such as Hall-MHD that neglects the electron inertia [19, 20] and extended MHD [21, 22, 23].
+The impact of pressure anisotropy was examined in [24, 25, 26, 27, 28].
+In addition, to include fast particle populations the problem was addressed in the frame of hybrid fluid-kinetic modes [29, 30, 31, 32, 33].
+For incompressible MHD flows having a poloidal component the generalized GS equation becomes elliptic (Eq.
+(3)) [34, 35] and decouples from the Bernoulli equation, which can be employed as a formula for the pressure (Eq.
+8).
+Incompressibility, implying that the density becomes uniform on the magnetic surfaces, is an acceptable approximation for fusion laboratory plasmas for two reasons: First, the poloidal velocities of those plasmas lie well within the first elliptic region.
+Second, since for typical flows in laboratory fusion plasmas, i.e.
+for Alfvén Mach numbers on the order of 0.01, density variations on the magnetic surfaces are very low.
+Equation (3), which contains five free surface quantities, has an additional electric-field term through the electrostatic potential $\Phi$, associated with the component of plasma velocity non-parallel to the magnetic field.
+For parallel velocity and arbitrary assignment of the free surface-function terms, an extension of the HELENA code was developed in [36], while the generalized Solov’ev solution to (3) (given by (12)), which constitutes a basic ingredient of the present study, was obtained in [35].
+Owing to the electric field, the respective equilibrium in addition to a separatrix, similar to that of the usual static Solov’ev equilibrium, possess a third X-point outside the separatrix (cf.
 Fig.
 11 of [35]).
-Other analytic solutions to the generalized GS equation of tokamak or space plasma pertinence were obtained by adopting a Solovév-like linearizing ansatz and imposing a diverted boundary [16]; alternative linearizing ansatzes [17, 18, 19]; non-linear forms of the generalized GS equation [20]; employing Lie-point symmetries for linear and nonlinear choices of the free-function terms [21, 22, 23]; and by the method of similarity reduction [24, 25].
+Other analytic solutions to the generalized GS equation of tokamak or space plasma pertinence were obtained by adopting a Solov’ev-like linearizing ansatz and imposing a diverted boundary [37]; alternative linearizing ansatzes [38, 39, 40]; non-linear forms of the generalized GS equation [41]; employing Lie-point symmetries for linear and nonlinear choices of the free-function terms [42, 43, 44]; and by the method of similarity reduction [45, 46].
 Aim of the present work is to construct analytically solutions to the generalized GS equation in its generic linearized form by employing an alternative method.
-This consists in pursuing solutions as expanded generalized Solovév ones, i.e.
-as superposition of the generalized Solovév solution and a function to be determined.
-In addition to the novelty of this method we are particularly interested in examining how the Solovév configuration is modified when the influence of the expanding function becomes stronger by changing the pertinent new free parameters.
-In section 2 the generalized GS equation and the generalized Solovév solution are briefly reviewed and the method of solving the generalized GS equation adopting the generic linearized ansatz is presented.
+This consists in pursuing solutions as expanded generalized Solov’ev ones, i.e.
+as superposition of the generalized Solov’ev solution and a function to be determined.
+In addition to the novelty of this method we are particularly interested in examining how the Solov’ev configuration is modified when the influence of the expanding function becomes stronger by changing the pertinent new free parameters.
+In section 2 the generalized GS equation and the generalized Solov’ev solution are briefly reviewed and the method of solving the generalized GS equation adopting the generic linearized ansatz is presented.
 Then, analytic solutions to the generic linearized generalized GS equation are constructed in section 3.
-In section 4 we examine how the Solovév configuration is modified by varying the additional free parameters associated with the pressure and the poloidal current in conjunction with the parallel component of the velocity and electric field, thus obtaining a variety of new configurations.
+In section 4 we examine how the Solov’ev configuration is modified by varying the additional free parameters associated with the pressure, the poloidal current in conjunction with the parallel component of the velocity, and electric field, thus obtaining a variety of new configurations.
 The conclusions are summarized in section 5.
@@
 The generalized GS equation is given in SI units by (cf. [34, 35])
+$(1-M_{p}^{2})\Delta^{\star}\psi-\frac{1}{2}\frac{dM_{p}^{2}}{d\psi}|\nabla\psi|^{2}+\frac{1}{2}\frac{d}{d\psi}\left(\frac{X^{2}}{1-M_{p}^{2}}\right)+\mu_{0}R^{2}\frac{dP_{s}}{d\psi}+\mu_{0}\frac{R^{4}}{2}\frac{d}{d\psi}\left[\frac{\rho}{1-M_{p}^{2}}\left(\frac{d\Phi}{d\psi}\right)^{2}\right]=0\,.$ (1)
+Here, $(z,R,\phi)$ are cylindrical coordinates with $z$ corresponding to the axis of symmetry; $\psi(R,z)$ is the poloidal magnetic flux function which labels the magnetic surfaces; $M_{p}(\psi)$ is the Mach function of the poloidal velocity with respect to the poloidal-magnetic-field Alfvén velocity; $\rho(\psi)$ and $\Phi(\psi)$ are the density and the electrostatic potential; $X(\psi)$ relates to the toroidal magnetic field and to the poloidal current (cf.
+relations (4) and (7)); for vanishing flow the surface function $P_{s}(\psi)$ coincides with the pressure (cf.
+(8)); $\Delta^{\star}=R^{2}\nabla\cdot(\nabla/R^{2})$ and $\mu_{0}$ is the permeability of free space.
+Equation (1) can be simplified by the transformation
+[…]
+which reduces (1) to
 $\Delta^{*}u+\frac{1}{2}\frac{d}{du}\left[\frac{X^{2}}{1-M_{p}^{2}}\right]+\mu_{0}R^{2}\frac{dP_{s}}{du}+\mu_{0}\frac{R^{4}}{2}\frac{d}{du}\left[\rho\left(\frac{d\Phi}{du}\right)^{2}\right]=0\,.$ (3)
-Here $(R,\phi,z)$ are cylindrical coordinate with $z$ corresponding to the axis of symmetry; the poloidal magnetic flux function $u(R,z)$ labels the magnetic surfaces; $M_{p}(u)$ is the Mach function of the poloidal velocity with respect to the poloidal Alfven velocity; $\Phi(u)$ is the electrostatic potential; $\rho(u)$ is the plasma density; $X(u)$ relates to the toroidal magnetic field $B_{\phi}=I/R$ through
-$I=\frac{X}{\left(1-M_{p}\right)}-R^{2}\frac{\sqrt{\rho}M_{p}\Phi^{\prime}}{\left(1-M_{p}\right)^{1/2}}\,.$ (2)
-The operator $\Delta^{*}$ in (1) is $\Delta^{*}:=\partial^{2}/\partial R^{2}-(1/R)\partial/\partial R+\partial^{2}/\partial z^{2}$ and $\mu_{0}$ is the permeability of free space.
-For vanishing flow, i.e.
-$M_{p}=\Phi^{\prime}=0$, the surface function $P_{s}(u)$ coincides with the pressure and Eq.
-(1) reduces to the GS equation.
-Following Ref. [14] we make the following choice of the free surface-function terms in (1)
+Note that no quadratic term as $|{\bf\nabla}u|^{2}$ appears anymore in (3).
+The transformation (2) does not affect the shape of magnetic surfaces; it just relabels them in terms of the function $u$.
+It is noted here that there is a misprint of the respective $R^{4}$-term in Eq.
+(17) of [35], i.e.
+the density factor should be linear instead of quadratic.
+To solve Eq.
+(3), the surface-quantity terms ($(1/2)(d/du)\left[X^{2}/(1-M_{p}^{2})\right]$, $\mu_{0}R^{2}dP_{s}/du$ and $\mu_{0}(R^{4}/2)(d/du)\left[\rho(d\Phi/du)^{2}\right]$) should be assigned as functions of $u$.
+Also, the equilibrium can be completely determined by calculations in the $u$-space; specifically, the magnetic field, velocity, current density, pressure and electric field are given by the relations:
+${\mathbf{B}}=I\nabla\phi+(1-M_{p}^{2})^{-1/2}\nabla\phi\times\nabla u\,,$ (4)
+with
+$I=(1-M_{p}^{2})^{-1/2}\left[\frac{X}{(1-M_{p}^{2})^{1/2}}-\mu_{0}R^{2}\sqrt{\frac{\rho}{\mu_{0}}}M_{p}\frac{d\Phi}{du}\right]\,,$ (5)
+${\mathbf{v}}=\frac{M_{p}}{\sqrt{\mu_{0}\rho}}{\mathbf{B}}-R^{2}(1-M_{p}^{2})^{-1/2}\frac{d\Phi}{du}\nabla\phi\,,$ (6)
+${\mathbf{j}}=\frac{1}{\mu_{0}}\left\{(1-M_{p})^{-1/2}\left[\Delta^{\star}u-\frac{1}{2(1-M_{p}^{2})}\frac{d(1-M_{p}^{2})}{du}|\nabla u|^{2}\right]\nabla\phi-\nabla\phi\times\nabla I\right\}\,,$ (7)
+$P=P_{s}-\rho\left[\frac{v^{2}}{2}-R^{2}\left(\frac{d\Phi}{du}\right)^{2}\right]\,,$ (8)
+${\mathbf{E}}=-\frac{d\Phi}{du}\nabla u$ (9)
+with $v^{2}={\mathbf{v}}\cdot{\mathbf{v}}$.
+Following Ref. [35] we first make the following choice of the free surface-function terms in (3)
 $\frac{d(\mu_{0}P_{s})}{du}$ $=$ $-\frac{P_{s0}}{U_{0}}\,,$
-$\frac{1}{2}\frac{d}{du}\left[\frac{\mu_{0}X^{2}}{1-M_{p}^{2}}\right]$ $=$ $\epsilon\frac{P_{s0}R_{0}^{2}}{U_{0}(1+\delta^{2})}\,,$
+$\frac{1}{2}\frac{d}{du}\left[\frac{X^{2}}{1-M_{p}^{2}}\right]$ $=$ $\epsilon\frac{P_{s0}R_{0}^{2}}{U_{0}(1+\delta^{2})}\,,$
 $\frac{1}{2}\frac{d}{du}\left[\mu_{0}\rho\left(\frac{d\Phi}{du}\right)^{2}\right]$ $=$ $-\lambda\frac{P_{s0}}{U_{0}(1+\delta^{2})R_{0}^{2}}\,.$ (10)
-Normalizing as $x:=R/R_{0}$, $y:=z/R_{0}$, $U:=u/u_{0}$, $u_{0}:=P_{s0}R_{0}^{4}/(2U_{0}(1+\delta^{2}))$, Eq. (1) assumes the form
-$U_{xx}-\frac{1}{x}U_{x}+U_{yy}-2\epsilon x^{4}+-2(1+\delta^{2})x^{2}+2\epsilon=0\,,$ (4)
-and the generalized Solovév solution is given by
+It is noted that a couple of the five surface quantities involved in the generalized GS equation remain free.
+Normalizing as $x:=R/R_{0}$, $y:=z/R_{0}$, $U:=u/u_{0}$ with $u_{0}:=P_{s0}R_{0}^{4}/(2U_{0}(1+\delta^{2}))$, Eq.
+(3) assumes the form
+$U_{xx}-\frac{1}{x}U_{x}+U_{yy}-2\lambda x^{4}-2(1+\delta^{2})x^{2}+2\epsilon=0\,,$ (11)
+and the generalized Solov’ev solution is given by
 $U_{gs}=\left[y^{2}\left(x^{2}-\epsilon\right)+\frac{\delta^{2}+\lambda}{4}(x^{2}-1)^{2}+\frac{\lambda}{12}(x^{2}-1)^{3}\right]\,.$ (12)
 The equilibrium configuration for $\lambda>0$ is shown in Fig. 1.
-Figure 1: The equilibrium configuration of the GS solution (5).
+Figure 1: The equilibrium configuration of the GS solution (12) on the poloidal plane $x-y$, with $x$ corresponds to the horizontal axis and $y$ to the vertical axis.
 The red-dashed curve represents the separatrix with $x_{s1}=\sqrt{\epsilon}$, $x_{s2}=\left(-\epsilon\lambda-3\delta^{2}+\sqrt{3}\sqrt{-\epsilon^{2}\lambda^{2}-2\delta^{2}\epsilon\lambda+3\delta^{4}+\delta^{2}\lambda+4\lambda^{2}}/(2\lambda)\right)^{1/2}$ and $y_{s}=\left[(1-\epsilon)^{2}(2\delta^{2}+\lambda(\epsilon+1))\right]^{1/2}$.
-As already mentioned in the Introduction, the outermost closed surface (red-dashed line) is a separatrix consisting of an elliptic outer part and a straight inner part parallel to the axis of symmetry; thus, a couple of X-points are created.
+As already mentioned in section 1, the outermost closed surface (red-dashed line) is a separatrix consisting of an elliptic outer part and a straight inner part parallel to the axis of symmetry; thus, a couple of X-points are created.
 Expressions of the X-point coordinates and of the inner and outer points of the configuration on the mid-plane $y=0$ in terms of the parameters $\delta$, $\epsilon$ and $\lambda$ are given in the caption of Fig.
@@
 For $\lambda=0$ the velocity becomes parallel to the magnetic field, while for $\lambda<0$ a third X-point appears outside the separatrix.
-When the velocity vanishes ($M_{p}=\lambda=0$), the usual static Solovév solution is recovered.
+When the velocity vanishes ($M_{p}=\lambda=0$), the usual static Solov’ev solution is recovered.
 Here we will consider equilibria with $\lambda>0$.
@@
 where $\bar{\alpha},\bar{\beta},\bar{\gamma}$ are dimensionless parameters.
-Setting $G:=2\epsilon$, $F:=-2(1+\delta^{2})$, $E:=-2\lambda$, $\alpha:=\bar{\alpha}(u_{0}/U_{0})$, $\beta:=\bar{\beta}(u_{0}/U_{0})$ and $\gamma:=\bar{\gamma}(u_{0}/U_{0})$ we obtain from (1) the following equation:
+The linear terms in $u$ in (2) result in more realistic equilibrium characteristics, e.g.
+the constant $\epsilon\left[P_{s0}R_{0}^{2}/(U_{0}(1+\delta^{2}))\right]$ in (2) results in monotonically increasing toroidal-current-density profiles while the additional linear in $u$ term, $-\bar{\beta}\epsilon\left[P_{s0}R_{0}^{2}/(U_{0}^{2}(1+\delta^{2})\right]u$, in (2) results in peaked toroidal-current-density profiles.
+Also, the linear term $\bar{\gamma}\left[\lambda P_{s0}/(U_{0}^{2}(1+\delta^{2}R_{0}^{2}))\right]u$ in (2) produces larger velocity shear.
+Setting $G:=2\epsilon$, $F:=-2(1+\delta^{2})$, $E:=-2\lambda$, $\alpha:=\bar{\alpha}(u_{0}/U_{0})$, $\beta:=\bar{\beta}(u_{0}/U_{0})$ and $\gamma:=\bar{\gamma}(u_{0}/U_{0})$ as well as substituting (2) into (3) we straightforwardly obtain the following equation:
 $U_{xx}-\frac{1}{x}U_{x}+U_{yy}+Ex^{4}+Fx^{2}+G-Ex^{4}\gamma U-Fx^{2}\alpha U-G\beta U=0\,.$ (14)
-Note that the solution to (7) for $\alpha=\beta=\gamma=0$ is given by (5). We now expand about the GS solution (5) as
+Note that the solution to (14) for $\alpha=\beta=\gamma=0$ is given by (12). We now will pursue a solution to (14) in the form of an expansion about the GS solution (12), i.e.
 $U=U_{gs}+W(x,y)\,,$ (15)
@@
 $=(\gamma Ex^{4}+\alpha Fx^{2}+\beta G)U_{gs}\,.$ (16)
-This is an inhomogeneous linear second-order partial differential equation for the unknown function $W(x,y)$.
+This is an inhomogeneous linear second-order PDE for the unknown function $W(x,y)$.
 The general solution consists of the sum of the general solution to the counterpart homogeneous equation plus a particular solution of the complete inhomogeneous equation.
@@
 where $k$ is the parameter of separation. The general solution of (19) is
-$Y(y)=D_{1}\cos(ky)+D_{2}\sin(ky)\,.$ (14)
+$Y(y)=C\cos(ky)+D\sin(ky)\,.$ (21)
 Setting $\xi=x^{2}$, Eq. (20) takes the form
@@
 $X_{2}(\xi)=cX_{1}(\xi)\ln\xi+\sum_{n=0}^{\infty}\bar{a}_{n}\xi^{n}\,,$ (26)
-where the coefficients $c$ and $\bar{a}_{n}$ can be determined by substituting (19) into (15). Here, we will employ the solution $X_{1}(\xi)$.
+where the coefficients $c$ and $\bar{a}_{n}$ can be determined by substituting (26) into (22).
+Here, we will employ the solution $X_{1}(\xi)$.
+By means of (18), (21), (23) and introducing a new parameter $s$ by $k=sl$ with $l=0,1,2,\ldots$, the solution of the homogeneous PDE (17) is written as:
+$W(x,y)=\sum^{\infty}_{l}X_{1}(\xi)\left[C_{l}\cos(sly)+D_{l}\sin(sly)\right]$
+$=\sum_{l=0}^{\infty}\sum_{n=0}^{\infty}a_{n,sl}\xi^{n}\left[C_{l}\cos(sly)+D_{l}\sin(sly)\right]\,.$ (27)
+The values of the free coefficients $C_{l}$ and $D_{l}$ may be properly fixed in order to impose a boundary.
+However, here since in the Solov’ev equilibrium the boundary (separatrix) is spontaneously formed we will not impose any boundary condition; accordingly, in (3) we will keep only the first two $l$-terms (for $l=0$ and $1$).
 Furthermore, we will pursue a particular solution of (2) in the form:
@@
 $W_{1}+\left(\xi-\frac{G}{2}\right):=\bar{W}_{1}\,,$ (32)
-Eq. (23) is put in the form:
+Equation (31) is put in the form:
 $4\xi\bar{W}_{1}^{\prime\prime}-(\gamma E\xi^{2}+\alpha F\xi+\beta G)\bar{W}_{1}=0\,.$ (33)
@@
 $\bar{W}_{1}=c_{1}Ai\left(\frac{\alpha F+\gamma E\xi}{(2\gamma E)^{2/3}}\right)+c_{2}Bi\left(\frac{\alpha F+\gamma E\xi}{(2\gamma E)^{2/3}}\right)\,,$ (34)
-where $Ai$ and $Bi$ are the Airy functions of the first and second kind. For $\beta\neq 0$ apart from the $k^{2}$-term, Eq. (15) is identical with (25). Therefore a solution of (25) is:
-$\bar{W}_{1}(\xi)=\xi\sum_{n=0}^{\infty}d_{n}\xi^{n}\,,$ (27)
-with the coefficients $d_{n}$ being determined by the recurrence relations (3) after setting $k=0$ (and replacing $a_{n}$ with $d_{n}$).
-Lastly, we will obtain a particular solution of the inhomogeneous equation (3). Introducing once more the variable $\xi=x^{2}$, Eq. (3) is written as:
+where $Ai$ and $Bi$ are the Airy functions of the first and second kind.
+For $\beta\neq 0$ apart from the $k^{2}$-term, Eq.
+(22) is identical to (33).
+Therefore the solution of (33) is already included in (3) for $l=0$.
+Finally, we will obtain a particular solution of the inhomogeneous equation (3). Introducing once more the variable $\xi=x^{2}$, Eq. (3) is written as:
 $4\xi W_{2}^{{}^{\prime\prime}}(\xi)-(\gamma E\xi^{2}+\alpha F\xi+\beta G)W_{2}=$
@@
 $C_{0}$ $=$ $\beta G\frac{(3F+2E+6)}{24}\,.$ (36)
-In view of the converging series solutions (18) and (27) to the respective homogeneous equations (15) and (25), we pursue a similar in form solution to (3):
+In view of the converging series solutions (25) to the homogeneous equations (22) and (33), we pursue a similarly structured solution to (3):
 $W_{2}=\xi\sum_{n=0}^{\infty}b_{n}\xi^{n}\,.$ (37)
@@
   1. 1.
-    When $\beta=0$, then $\epsilon=0$ and therefore the second of (2) becomes $(1/2)d/du(X^{2}/(1-M_{p}^{2}))=0$.
+    When $\beta=0$, then $\epsilon=0$ and therefore the second of (2) becomes
+    $(1/2)d/du(X^{2}/(1-M_{p}^{2}))=0$.
 In the absence of flow then (5) implies that the plasma is confined by a vacuum toroidal magnetic field ($\beta_{p}=1$ equilibrium); when the non-parallel flow is there the electric-field term in (5) additionally contributes to the poloidal current.
@@
 In this case all the linear terms in the ansatz (2) can remain finite.
-The latter relation just acts as a constraint on the parameters $\epsilon$, $\delta$ and $\lambda$.
-To construct the equilibria of Sec.
-4 we treated $\epsilon$ and $\lambda$ as free and
+This relation just acts as a constraint on the parameters $\beta$, $\delta$ and $\lambda$.
+In section 4 we intend to examine the impact of the linear in $u$ electric-field term in (2) on the equilibrium by varying the parameter $\beta$; accordingly, leaving $\beta$ free we will employ the constraint to express $\delta$ in terms of $\beta$ and $\lambda$, i.e.
     $\delta=\left[\frac{(2/3)(6-\beta\lambda)}{\beta}\right]^{1/2}\,.$ (39)
@@
     When $\epsilon=0$, $\beta$ does not necessarily vanish and both $\delta$ and $\lambda$ remain free.
-Additional free parameters are $\alpha$, $\beta$, $\gamma$, $a_{0}$, $d_{0}$, $b_{0}$, $D_{1},D_{2}$, $R_{0}$, $P_{s0}$, $U_{0}$ and $k$.
+Additional free parameters are $\alpha$, $\beta$, $\gamma$, $a_{0}$, $b_{0}$, $C_{0}$, $C_{1}$, $D_{1}$, $R_{0}$, $P_{s0}$, $U_{0}$ and $s$.
+It is recalled here that $R_{0}$ is a reference length determining the $R$-position of the magnetic axis; the parameters $P_{s0}$ and $U_{0}$ are reference parameters associated with the pressure and poloidal magnetic flux; $\delta$ and $\epsilon$ are geometrical (shaping) parameters of the starting generalized Solov’ev equilibrium (12) shown in Fig.
+1; $\lambda$ relates to the electric field term in (12) through the ansatz (2); $\alpha$, $\beta$ and $\gamma$ are physical parameters of the current density, pressure and electric field, respectively, in connection to the additional linear terms in the ansatz (2); and $a_{0}$, $b_{0}$, $C_{0}$, $C_{1}$, $D_{1}$ and $s$ are shaping parameters associated with the new constructed solution.
+A list of the free parameters is given in Table 1.
 Also, it is noted that the series expansion (37) was shown numerically to converge and we kept terms up to the order of $\xi^{20}$ in our calculations.
-Summarizing, a solution of the general linearized form of the generalized GS equation (2) consists of a separated solution of its homogeneous counterpart in terms of (11), (14) and (18) plus a particular solution of the form (20), expressed in terms of the series (27) and (30).
-Figure 2: Left: different equilibrium boundaries are given as the additional linear-in-$U$, pressure term in the first of (2) varies in connection with different values of $\alpha$.
-The black continuous curve represents the separatrix of the respective genaralized Solovév spheromak configuration.
-Right: a set of magnetic surfaces for one of the configuration representd by blue, dot-dashed U-curves are shown together with the respective curves of the genaralized Solovév equilibrium.
+Table 1: List of the free parameters
+Parameters Description
+$R_{0}$ Reference $R$-position of the magnetic axis
+$P_{s0}$, $U_{0}$ Reference pressure and poloidal magnetic flux
+$\delta$, $\epsilon$ Shaping of the generalized Solov’ev equilibrium of Fig. (1)
+$\lambda$ Parameter related to the electric field term in the
+generalized Solov’ev solution (12)
+$\alpha$, $\beta$, $\gamma$ Current density, pressure and electric field parameters
+in the ansatz (2)
+$a_{0}$, $b_{0}$, $C_{0}$, $C_{1}$, $D_{1}$, $s$ Shaping parameters associated with the new constructed
+solution
+$\rho_{0}$, $l_{\rho}$, $M_{p0}$, $l_{M_{p}}$ Density and poloidal Mach function parameters (Eq. (40))
+In summary, a solution of the general linearized form of the generalized GS equation (2) consists of a separated solution of its homogeneous counterpart in terms of (18), (21) and (25) plus a particular solution of the form (28), expressed in terms of the series (3) for $l=0$ and (37).
+Figure 2: Left: different equilibrium boundaries are given on the poloidal plane $x-y$ as the additional linear-in-$U$ pressure term in the first of (2) varies in connection with different values of $\alpha$.
+The black continuous curve represents the separatrix of the respective genaralized Solov’ev spheromak configuration.
+Right: a set of magnetic surfaces for one of the configuration representd by blue, dot-dashed U-curves are shown together with the respective curves of the genaralized Solov’ev equilibrium.
+$U_{b}$ is the value of the bounding $u$-curve (separatrix).
 4 Equilibrium configurations
-By employing the solution to (7) derived in the previous section we have constructed several new equilibria.
-Particularly, we examined how the generalized Solovév equilibrium is modified by varying the additional parameters $\alpha$, $\beta$ and $\gamma$, associated with the respective linear-in-$u$ terms in (2) of pressure, poloidal-current in conjunction with the poloidal velocity, and electric field.
-To impose a stagnation point, $(x_{0},y_{0})$, in addition to the conditions $dU/dx|_{x_{0},y_{0}}$ and $dU/dy|_{x_{0},y_{0}}$ we considered the sign of the quantity
+By employing the solution to (14) derived in section 3 we have constructed several new equilibria.
+Particularly, we examined how the generalized Solov’ev equilibrium is modified by varying the additional parameters $\alpha$, $\beta$ and $\gamma$, associated with the respective linear-in-$u$ terms in (2) of pressure, poloidal-current in conjunction with the parallel velocity (Eqs.
+(5) and (7)), and electric field.
+To impose a stagnation point, $(x_{0},y_{0})$, in addition to the conditions $dU/dx|_{x_{0},y_{0}}=0$ and $dU/dy|_{x_{0},y_{0}}=0$ we considered the sign of the quantity
 $\left.\frac{d^{2}U}{dx^{2}}\frac{d^{2}U}{dy^{2}}-\left(\frac{d^{2}U}{dxdy}\right)^{2}\right|_{(x_{0},y_{0})}\,,$
 which should be positive for a magnetic axis and negative for an X-point.
-For the up-down symmetric configurations considered, the magnetic axis is located at the same position, $(x=1,y=0)$ as the generalized Solovév configuration and $U=U_{a}=0$ thereon.
-Note that in this case the condition $dU/dy|_{(1,0)}$ is identically satisfied because of symmetry.
-If not otherwise specified, we assigned the following parametric values: $k=0.1$, $\delta=0.1$, $\lambda=0.1$, $D_{1}=1$ and $D_{2}=0$.
+For the up-down symmetric configurations considered, the magnetic axis is located at the same position, $(x=1,y=0)$ as the generalized Solov’ev configuration and $U=0$ thereon.
+Note that in this case the condition $dU/dy|_{(1,0)}=0$ is identically satisfied because of symmetry.
+If not otherwise specified, we assigned the following parametric values: $s=0.1$, $\delta=0.1$, $\lambda=0.1$ and $a_{0}=1$.
 It is recalled that when $\beta\neq 0$ and $\epsilon\neq 0$, $\delta$ should be calculated from (39).
-The parameters $a_{0}$, $b_{0}$ and $d_{0}$ were determined as solutions to the set of the algebraic equations derived by the conditions associated with the imposition of a stagnation point.
+The parameters $C_{0}$, $C_{1}$ and $D_{1}$ were determined as solutions to the set of the algebraic equations derived by the conditions associated with the imposition of a stagnation point.
 Since we employed dimensionless quantities, the reference dimensional parameters $R_{0}$, $P_{s0}$ (in units of $4\pi\times 10^{-7}$ H/mPA ) and $U_{0}$ remained free.
 Depending on the value of $\epsilon$ we constructed tokamak, spherical tokamak and spheromak equilibria, the latter corresponding to $\epsilon=0$.
-First, we examined the impact of the $\alpha$- $\beta$- and $\gamma$-term individually on up-down symmetric equilibria equilibria by imposing the same position of the magnetic axis, $(x=1,y=0)$, as that of the respective generalized Solovév equilibrium and the same value $U=0$ thereon.
+First, we examined the impact of the $\alpha$- $\beta$- and $\gamma$-terms individually on up-down symmetric equilibria by imposing the same position of the magnetic axis, $(x=1,y=0)$, as that of the respective generalized Solov’ev equilibrium and the same value $U=0$ thereon.
 In the presence of $\alpha$-term, a D-shaped configuration is formed with negative triangularity, the boundary of which is shown in Fig.
 2-left by the red-dashed curve.
-As $\alpha$ gradually takes larger values the configuration increasing triangularity extends towards the axis of symmetry, touches the axis of symmetry forming a separatrix with a single X-point located on the origin of the coordinate system (blue, dot-dashed line) and then becomes a peculiar configuration having a separatrix with a couple of X-points (green-dotted line).
+As $\alpha$ gradually takes larger values the configuration with increasing triangularity extends toward the axis of symmetry, eventually touching it to form a separatrix with a single X-point located at the origin of the coordinate system (blue, dot-dashed line) and then becomes a peculiar configuration having a separatrix with a couple of X-points (green-dotted line).
 The $\beta$-term creates configurations with small triangularity elongated along the $x$-axis as shown in Fig.
 3.
-The red-dashed curves represent the magnetic surfaces of a configuration with the same poloidal magnetic flux as the respective generalized Solovév tokamak configuration represented by the black-continuous curve.
-As $\beta$ takes larger values the configuration gets more extended and eventually reaches the axis of symmetry becoming a spheromak one.
+The red-dashed curves represent the magnetic surfaces of a configuration with the same poloidal magnetic flux as the respective generalized Solov’ev tokamak configuration represented by the black-continuous curve.
+As $\beta$ takes larger values the configuration extends further and eventually reaches the axis of symmetry transforming into a spheromak.
 It is noted that for $\epsilon=\beta=\gamma=0$ the equilibrium becomes independent of $\beta$.
@@
 In Fig.
-5 the blue, dot-dashed curves represent an up-down symmetric spherical tokamak configuration ($\epsilon=0.1$) the X-points of which are common with those of the respective generalized Solovév equilibrium, but the poloidal magnetic flux of this former equilibrium is larger than that of the latter.
-The red-dashed curve represents the boundary of a D-shaped configuration with poloidal magnetic flux equal to that of the generalized Solovév solution.
+5 the blue, dot-dashed curves represent an up-down symmetric spherical tokamak configuration ($\epsilon=0.1$) the X-points of which are common with those of the respective generalized Solov’ev equilibrium, but the poloidal magnetic flux of this former equilibrium is larger than that of the latter.
+The red-dashed curve represents the boundary of a D-shaped configuration with poloidal magnetic flux equal to that of the generalized Solov’ev solution.
 In Fig.
@@
 7 shows an up-down asymmetric configuration with the magnetic axis located at ($x=1.02,y=0.83$) and the single lower X-point at $(x=0.94,y=-22.6)$.
-Figure 3: The impact of the additional linear-in-$U$, poloidal-current term in the second of (2) in connection with different values of $\beta$.
-The black continuous curves represent the magnetic surfaces of the respective generalized Solovév tokamak configuration.
+Figure 3: Fig.
+3 shows the impact of the additional linear-in-$U$, poloidal-current term in the second of (2) in connection with different values of $\beta$.
+The black continuous curves represent the magnetic surfaces of the respective generalized Solov’ev tokamak configuration.
+$U_{b}$ is the value of the bounding $u$-curve (separatrix).
 Figure 4: Left: different equilibrium boundaries are given as the additional linear-in-$U$, electric-field term in the third of (2) varies in connection with different values of $\gamma$.
-The black continuous curve represents the separatrix of the respective generalized Solovév spheromak configuration.
-Right: A set of magnetic surfaces for one of the configuration represented by green-dotted U-curves are shown together with the respective curves of the Solovév equilibrium.
-Figure 5: A spherical tokamak diverted configuration ($\epsilon=0.1$, $\delta=20$) with a couple of X-points common with the respective Solovév equilibrium.
-The red-dashed curve represents the boundary of a D-shaped tokamak equilibrium with poloidal magnetic flux equal to that of the Solovév one.
+The black continuous curve represents the separatrix of the respective generalized Solov’ev spheromak configuration.
+Right: A set of magnetic surfaces for one of the configuration represented by green-dotted U-curves are shown together with the respective curves of the Solov’ev equilibrium.
+$U_{b}$ is the value of the bounding $u$-curve (separatrix).
+Figure 5: A spherical tokamak diverted configuration ($\epsilon=0.1$, $\delta=20$) with a couple of X-points common with the respective Solov’ev equilibrium.
+The red-dashed curve represents the boundary of a D-shaped tokamak equilibrium with poloidal magnetic flux equal to that of the Solov’ev one.
+$U_{b}$ is the value of the bounding $u$-curve (separatrix).
 Figure 6: An up-down symmetric, diverted tokamak configuration ($\epsilon=0.1,\ \delta=20$) with a couple of X-points located at ($x=0.84,y=\pm 14.5$).
-The black continuous curves represent the magnetic surfaces of the respective Solovév equilibrium.
+The black continuous curves represent the magnetic surfaces of the respective Solov’ev equilibrium.
+$U_{b}$ is the value of the bounding $u$-curve (separatrix).
 Figure 7: An up-down asymmetric, diverted tokamak configuration for $k=0.2$, $\epsilon=0.2$, $\delta=31.6$, $D_{1}=0.5$ and $D_{2}=0.01$.
 The magnetic axis is located at $(x=1.02,y=0.83)$ and the single lower X-point at $(x=0.94,y=-22.6)$.
-The black continuous curves represent the magnetic surfaces of the respective Solovév equilibrium.
+The black continuous curves represent the magnetic surfaces of the respective Solov’ev equilibrium.
+$U_{b}$ is the value of the bounding $U$-curve (separatrix).
+Finally, to completely determine the equilibrium two from the five free functions involved should be specified.
+For example, considering as free functions the density and the poloidal Mach function we make the choice
+$\rho=\rho_{0}\left(1-\frac{U}{U_{b}}\right)^{l_{\rho}}\,,\ \ M_{p}=M_{p0}\left(1-\frac{U}{U_{b}}\right)^{l_{M_{p}}}\,,$ (40)
+where $\rho_{0}$, $M_{p0}$, $l_{\rho}$ and $l_{M_{p}}$ are free parameters. Introducing the dimensionless quantities
+$\tilde{\rho}=\frac{\rho}{\rho_{0}}\,,\ \ \tilde{X}=\frac{X}{u_{0}/R_{0}}\,,\ \ \tilde{{\mathbf{B}}}=\frac{{\mathbf{B}}}{u_{0}/R_{0}^{2}}\,,\ \ \tilde{{\mathbf{j}}}=\frac{{\mathbf{j}}}{u_{0}/(R_{0}^{3}\mu_{0})}\,,\ \ \tilde{{\mathbf{v}}}=\frac{{\mathbf{v}}}{u_{0}/(R_{0}^{2}\sqrt{\mu_{0}\rho_{0}})}\,,$
+$\tilde{\Phi}=\frac{\Phi}{u_{0}^{2}/(R_{0}^{3}\sqrt{\mu_{0}\rho_{0}})}\,,\ \ \tilde{P}=\frac{P}{P_{s0}/(u_{0}^{2}/R_{0}^{4})}\,,$
+integration of the relations (2) with respect to $u$ yields (in the dimensionless form and omitting the tildes)
+$P_{s}$ $=$ $A-2(1+\delta^{2})U+\alpha(1+\delta^{2})U^{2}$
+$\frac{X^{2}}{1-M_{p}^{2}}$ $=$ $B+4\epsilon U-2\beta U^{2}$
+$\rho\left(\frac{d\Phi}{dU}\right)^{2}$ $=$ $C-4\lambda U+2\gamma U^{2}\,,$ (41)
+where $A$, $B$ and $C$ are free parameters. The reference density $\rho_{0}$ remains free. Using (4) in (8) we obtain
+$P=A-2(1+\delta^{2})U+\alpha(1+\delta^{2})U^{2}-\frac{1}{2}\rho v^{2}+x^{2}\left(C-4\lambda U+2\gamma U^{2}\right).$ (42)
+The values of $A$ and $C$ are fixed so that the pressure vanish on the boundary.
+Isobaric contours for the D-shaped tokamak configuration of Fig.
+5, the boundary of which is indicated by the red-dotted curve therein, are given in Fig.
+8 together with pressure profiles on the poloidal plane at $y=0$ and at the vertical line passing from the position of the pressure maximum.
+It is noted that although the values of the dimensionless pressure are quite high, realistic pressure values can be obtained by appropriate values of the free parameters $P_{s0}$ and $U_{0}$.
+For flows of experimental fusion pertinence the deviation of the magnetic surfaces from the isobaric surfaces is very small.
+Figure 8: Upper: Isobaric (blue-continuous) curves on the poloidal plane for the D-shaped tokamak configuration of Fig.
+(5) for $B=0.2$, $M_{p0}=0.1$, $l_{\rho}=2$ and $l_{M_{p}}=2$.
+The red-dashed curves represent magnetic surfaces.
+Lower left: Respective pressure profile on the mid-plane $y=0$.
+Lower right: Respective vertical pressure profile at the position of the pressure maximum
 5 Conclusions
 Making a generic linearizing choice of the free function-terms involved in the generalized GS equation (3) we solved the resulting linearized equation (2) analytically.
-The requested solution was expressed as an expansion of the generalized Solovév solution (5) in the form (8) involving the determinable function $W$.
+The requested solution was expressed as an expansion of the generalized Solov’ev solution (12) in the form (15) involving the determinable function $W$.
 Then, the solution was obtained as a superposition of the counterpart homogeneous equation and a particular solution of the complete inhomogeneous equation in the form (28).
 The solutions of the pertinent radial ODEs were expressed in terms of infinite converging series.
-Employing these solutions, we constructed several configurations by changing the values of the free parameters, either individually or in combination, associated with the new terms of pressure, poloidal current and electric field.
+Employing these solutions, we constructed several configurations by changing the values of the free parameters, either individually or in combination, associated with the new terms of pressure, poloidal current in conjunction with the parallel velocity, and electric field.
 The equilbria constructed are pertinent to tokamaks, spherical tokamaks and spheromaks; they include D-shaped configurations with either positive or negative triangularity and diverted configurations with either a couple of X-points or a single X-point.
@@
 This work was conducted in the framework of participation of the University of Ioannina in the National Programme for the Controlled Thermonuclear Fusion, Hellenic Republic.
+The authors would like to thank the anonymous reviewers for constructive comments, which have resulted in a remarkably improved version of the manuscript.